# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ 2.5 Efficient influence functions of the target causal parameters
For our proposed estimator of $\Psi:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$, we use the adaptive-TMLE for $\Psi^{\#}$ that first learns a parametric working model ${\cal T}_{n}=\{\tau_{w,n,\beta}:\beta\}$ for $\tau_{0}$ that will approximate the true bias function $\tau_{P_{0}}$, implying a corresponding semiparametric regression working model ${\cal M}_{w,2,n}$ for $P_{0}$, and then constructs an efficient estimator for the corresponding projection parameter $\Psi^{\#}(\Pi_{{\cal M}_{w,2,n}}(P_{0}))$, where $\Pi_{\mathcal{M}_{w,2,n}}(P_{0})$ is a projection of $P_{0}$ onto the working model $\mathcal{M}_{w,2,n}$ [adml].
@@
4.1 Semiparametric regression working model for the conditional effect of the study indicator
@@
Since $\bar{Q}_{0}(S,W,A)=\bar{Q}_{0}(0,W,A)+S\tau_{0}(W,A)$, a working model ${\cal T}_{n}$ for $\tau_{0}$ corresponds with a semiparametric regression working model ${\cal Q}_{w,n}=\{\theta+S\tau_{w,n,\beta}:\beta,\theta\}$ for $\bar{Q}_{0}=E_{0}(Y\mid S,W,A)$, where $\theta$ is an unspecified function of $W,A$. The working model ${\cal Q}_{w,n}$ further implies a working model ${\cal M}_{w,2,n}=\{P\in{\cal M}:\bar{Q}_{P}\in{\cal Q}_{w,n}\}\subset{\cal M}$ for the observed data distribution $P_{0}$. Let $\tau_{w,n,\beta(P)}\in{\cal T}_{n}$ be a projection of $\tau_{P}$ onto this working model ${\cal T}_{n}$ with respect to some loss function. We will also use the notation $\tau_{w,n,P}$ for $\tau_{w,n,\beta(P)}$. This generally corresponds with mapping a $P\in{\cal M}$ into a projection $\Pi_{w,n}(P)\in{\cal M}_{w,2,n}$, and setting $\tau_{w,n,P}\equiv\tau_{\Pi_{w,n}(P)}$. For the squared-error loss, and a semiparametric regression working model ${\cal Q}_{w,n}$ for $\bar{Q}_{P}$, we recommend the projection used in [adml]:
$\bar{Q}_{w,n,P}\equiv\Pi_{{\cal Q}_{w,n},P}(\bar{Q}_{P})=\arg\min_{\bar{Q}\in{\cal Q}_{w,n}}P\left(\bar{Q}_{P}-\bar{Q}\right)^{2}.$
Interestingly, as shown in [adml], we have
$\tau_{w,n,P}=\arg\min_{\tau\in\tau_{n}}E_{P}L_{\tilde{\theta}_{P},\Pi_{P}}(\tau),$
@@
$L_{\tilde{\theta}_{P},\Pi_{P}}(\tau)\equiv\left(Y-\tilde{\theta}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))\tau\right)^{2},$
and $\tilde{\theta}_{P}(W,A)=E_{P}(Y\mid W,A)$. This insight provides us with a nice loss function for learning a working model ${\cal T}_{n}$, by using, for example, a highly adaptive lasso minimum-loss estimator (HAL-MLE) [undersmooth] for this loss function with estimators $\tilde{\theta}_{n}$ of $\tilde{\theta}_{0}$ and $\Pi_{n}$ of $\Pi_{0}$. Specifically, given a rich linear model $\{\tau_{\beta}=\sum_{j}\beta_{j}\phi_{j}\}$ with a large set of spline basis functions $\{\phi_{j}:j\}$, we compute the lasso-estimator
$\beta_{n}=\arg\min_{\beta,\parallel\beta\parallel_{1}\leq C_{n}}\sum_{i=1}^{n}\left(Y_{i}-\tilde{\theta}_{n}(W_{i},A_{i})-(S_{i}-\Pi_{n}(1\mid W_{i},A_{i}))m_{\beta}(W_{i},A_{i})\right)^{2},$
where $m_{\beta}(W,A)$ is a linear combination of HAL basis functions and $C_{n}$ is an upper bound on the sectional variation norm [benkeser_hal_2016]. The working model for $\tau_{P}$ is then given by ${\cal T}_{n}=\{\sum_{j,\beta_{n}(j)\not=0}\beta_{j}\phi_{j}:\beta\}$, i.e. linear combinations of the basis functions with non-zero coefficients. In addition, it has been shown in [adml] that
$\tau_{w,n,P}=\arg\min_{\tau\in\mathcal{T}_{n}}E_{P}\Pi_{P}(1-\Pi_{P})(1\mid A,W)(\tau_{P}-\tau)^{2}.$
Put in other words, the squared-error projection of $\bar{Q}_{P}$ onto the semi-parametric regression working model corresponds with a weighted squared-error projection of $\tau_{P}$ onto the corresponding working model for $\tau_{P}$. The weights $\Pi_{P}(1-\Pi_{P})(1\mid W,A)$ stabilize the efficient influence function for $\beta(P)$, showing that the loss function $L_{\tilde{\theta},\Pi}(\tau)$ for $\tau_{P}$ yields relatively robust estimators of $\tau_{0}$.
4.2 Data-adaptive working-model-specific projection parameter for $\Psi^{\#}$
Given the above definition of the semiparametric regression working model, $\tau_{w,n,P}=\tau_{w,n,\beta(P)}$ for $P\in\mathcal{M}$, for the conditional effect of the study indicator $S$ on the outcome $Y$, we can now define a corresponding working-model-specific projection parameter for $\Psi^{\#}(P)$. Specifically, let $\Psi^{\#}_{{\cal M}_{w,2,n}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ be defined as
$\Psi^{\#}_{{\cal M}_{w,2,n}}(P)\equiv E_{P}\Pi_{P}(0\mid W,0)\tau_{w,n,\beta(P)}(W,0)-E_{P}\Pi_{P}(0\mid W,1)\tau_{w,n,\beta(P)}(W,1).$
We will sometimes use the notation $\Psi^{\#}(P_{W},\Pi_{P},\tau_{w,n,P})$ to emphasize its reliance on the nuisance parameters $(P_{W},\Pi_{P},\tau_{w,n,P})$. The following lemma reviews the above stated results.
Lemma 3.
Let ${\cal Q}_{w}=\{\theta(A,W)+ST_{\beta}(W,A):\theta,\beta\}$, where $T_{\beta}(W,A)=\sum_{j}\beta_{j}\phi_{j}(W,A)$. Let $\bar{Q}_{w,P}=\arg\min_{\bar{Q}\in{\cal Q}_{w}}P(\bar{Q}_{P}(S,W,A)-\bar{Q}(S,W,A))^{2}$, where $\bar{Q}_{P}=E_{P}(Y\mid S,W,A)$. Let $\tau_{P}(W,A)=E_{P}(Y\mid S=1,W,A)-E_{P}(Y\mid S=0,W,A)$. We have that $\bar{Q}_{w,P}=E_{P}(Y\mid W,A)+\sum_{j}(S-\Pi_{P}(1\mid W,A))\beta_{P}(j)\phi_{j}(W,A)$, where
$\beta(P)=\arg\min_{\beta}E_{P}\Pi_{P}(1-\Pi_{P})(1\mid W,A)\left(\tau_{P}(W,A)-\sum_{j}\beta_{j}\phi_{j}(W,A)\right)^{2}.$
Thus, $\sum_{j}\beta_{P}(j)\phi_{j}$ represents the projection of $\tau_{P}(W,A)$ onto $\{\sum_{j}\beta_{j}\phi_{j}:\beta\}$ with respect to a weighted $L^{2}$-norm with weights $\Pi_{P}(1-\Pi_{P})(1\mid W,A)$. We also have
$\beta(P)=\arg\min_{\beta}E_{P}\left(\bar{Q}_{P}-(S-\Pi_{P}(1\mid W,A))\sum_{j}\beta_{j}\phi_{j}(W,A)\right)^{2},$
and thereby
$\beta(P)=\arg\min_{\beta}E_{P}\left(Y-(S-\Pi_{P}(1\mid W,A))\sum_{j}\beta_{j}\phi_{j}(W,A)\right)^{2}.$
Here, one could replace $Y$ by $Y-E_{P}(Y\mid W,A)$ as well.
This also shows that
$T_{P}=\arg\min_{T}E_{P}(Y-E_{P}(Y\mid W,A)-(S-\Pi_{P}(1\mid W,A))T(W,A))^{2}.$
Therefore, we can use the following loss function for learning $T_{P}$:
$L_{\tilde{\theta}_{P},\Pi_{P}}(T)=(Y-\tilde{\theta}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))T(W,A))^{2},$
where the loss function is indexed by two nuisance parameters $\tilde{\theta}_{P}(W,A)=E_{P}(Y\mid W,A)$ and $\Pi_{P}(1\mid W,A)=P(S=1\mid W,A)$.
4.3 Canonical gradient of $\Psi^{\#}_{{\cal M}_{w,2}}$ at $P$
Let’s first find the canonical gradient of the $\beta(P)$-component.
Lemma 4.
Let $I_{P}=E_{P}\Pi(1-\Pi)(1\mid W,A){\bf\phi}{\bf\phi}^{\top}(W,A)$. Let $\beta(P)$ be defined as in Lemma 3 on a nonparametric model. The canonical gradient of $\beta$ at $P$ is given by:
$D_{\beta,P}=I_{P}^{-1}(S-\Pi_{P}(1\mid W,A)){\bf\phi}(W,A)\left(Y-\tilde{\theta}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))\sum_{j}\beta_{P}(j)\phi_{j}(W,A)\right),$
where $\tilde{\theta}_{P}=E_{P}(Y\mid W,A)$, $\Pi_{P}(1\mid W,A)=P(S=1\mid W,A)$. This can also be written as
$D_{\beta,P}=I_{P}^{-1}(S-\Pi_{P}(1\mid W,A)){\bf\phi}(W,A)(Y-\bar{Q}_{w,P}(S,W,A)).$
The proof can be found in Appendix A. We then need to find the canonical gradient of the working-model-specific projection parameter $\Psi^{\#}_{{\cal M}_{w,2}}$ at $P$.
Lemma 5.
The canonical gradient of $\Psi^{\#}_{{\cal M}_{w,n}}$ at $P$ is given by
$D^{\#}_{{\cal M}_{w,2},P}=D^{\#}_{{\cal M}_{w,2},W,P}+D^{\#}_{{\cal M}_{w,2},\Pi,P}+D^{\#}_{{\cal M}_{w,2},\beta,P},$
@@ 4.3 Canonical gradient of $\Psi^{\#}_{{\cal M}_{w,2}}$ at $P$
$D^{\#}_{{\cal M}_{w,2},W,P}$ $\equiv\Pi_{P}(0\mid W,0)\sum_{j}\beta_{P}(j)\phi_{j}(W,0)-\Pi_{P}(0\mid W,1)\sum_{j}\beta_{P}(j)\phi_{j}(W,1)-\Psi^{\#}_{{\cal M}_{w,2}}(P)$
$D^{\#}_{{\cal M}_{w,2},\Pi,P}$ =$=\left\{\frac{A}{g_{P}(1\mid W)}\tau_{w,\beta(P)}(W,1)-\frac{1-A}{g_{P}(0\mid W)}\tau_{w,\beta(P)}(W,0)\right\}(S-\Pi_{P}(1\mid W,A))$
$D^{\#}_{{\cal M}_{w,2},\beta,P}$ $=\sum_{j}D_{\beta,P,j}E_{P}\Pi_{P}(0\mid W,0)\phi_{j}(W,0)-\sum_{j}D_{\beta,P,j}E_{P}\Pi_{P}(0\mid W,1)\phi_{j}(W,1),$
where
$D_{\beta,P}=I_{P}^{-1}(S-\Pi_{P}(1\mid W,A)){\bf\phi}(W,A)(Y-\bar{Q}_{w,P}(S,W,A)).$
@@
4.4 TMLE of the projection parameter $\Psi^{\#}_{{\cal M}_{w,2,n}}$
@@
A plug-in estimator of the projection parameter $\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{0})$ requires an estimator of $\Pi_{0}$, $\beta_{0}$, $P_{W,0}$, and a model selection method, as discussed earlier, so that the semiparametric regression working model ${\cal M}_{w,2,n}$ is known, including the working model $\{\tau_{w,n,\beta}:\beta\}$ for $\tau_{P_{0}}$. We then need to solve each component of the canonical gradient of $\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{0})$. First, for the $W$-component, we use the empirical distribution $P_{W,n}$ of $W$ as the estimator of $P_{W,0}$. Therefore, $P_{n}D^{\#}_{\mathcal{M}_{w,2,n},W,\beta,\Pi,P_{W,n}}(O)=0$ for any $\beta,\Pi$. Then, for the $\beta$-component, we need to construct an initial estimator $\tilde{\theta}_{n}$ of $\tilde{\theta}_{0}(W,A)=E_{0}(Y\mid W,A)$ and $\Pi_{n}$ of $\Pi_{0}(s\mid W,A)=P_{0}(S=s\mid W,A)$. Then, with these two nuisance estimates, we can compute the MLE over the working model for $\tau_{0}$ with respect to our loss function $L_{\tilde{\theta}_{n},\Pi_{n}}(\tau)$:
$\beta_{n}=\arg\min_{\beta}P_{n}L_{\tilde{\theta}_{n},\Pi_{n}}(\tau_{w,n,\beta}).$
@@
We have then solved $P_{n}D^{\#}_{{\cal M}_{w,2,n},\beta,\beta_{n},\Pi_{n}}(O)=0$. Finally, for the $\Pi$-component, note that we have the representation: $D^{\#}_{{\cal M}_{w,2,n},\Pi,P}(O)=C(g,\beta)(S-\Pi_{P}(1\mid W,A))$, where we refer to $C(g,\beta)$ as the clever covariate. With an estimator $g_{n}$ of $g_{0}$, we can compute a targeted estimator $\Pi_{n}^{*}$ solving $P_{n}C(g_{n},\beta_{n})(S-\Pi_{n}^{*}(1\mid W,A))=0$. The desired TMLE is then given by $\Psi^{\#}_{\mathcal{M}_{w,2,n}}(P_{n}^{*})\equiv\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{W,n},\Pi_{n}^{*},\beta_{n})$. For inference with respect to the projection parameter $\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{0})$, we use that
$\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{n}^{*})-\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{0})=P_{n}D^{\#}_{{\cal M}_{w,2,n},P_{0}}+o_{P}(n^{-1/2}),$

### v2

@@ 4.3 TMLE of the projection parameter $\tilde{\Psi}_{{\cal M}_{A,w}}$
We now discuss the construction of an A-TMLE for the bias estimand, $\Psi^{\#}(P_{0})=E_{0}[\Pi_{0}(0\mid W,0)\tau_{S,0}(W,0)-\Pi_{0}(0\mid W,1)\tau_{S,0}(W,1)]$. The steps involve first data-adaptively learning a working model for the conditional effect of the study indicator $S$ on the outcome $Y$. That is, a working model $\mathcal{T}_{S,w}=\{\tau_{S,\beta}:\beta\}$ for $\tau_{S,0}$. This working model implies a semiparametric regression working model $\mathcal{M}_{S,w}$ for $P_{0}$. We then construct an efficient estimator for the corresponding projection estimand $\Psi^{\#}_{\mathcal{M}_{S,w}}(P_{0})\equiv\Psi^{\#}(\Pi_{\mathcal{M}_{S,w}}(P_{0}))$.
@@
5.1 Semiparametric regression working model for the conditional average RCT-enrollment effect
@@
Since $Q_{0}(S,W,A)=Q_{0}(0,W,A)+S\tau_{S,0}(W,A)$, a working model $\mathcal{T}_{S,w}$ for $\tau_{S,0}$ corresponds to a semiparametric regression working model ${\cal Q}_{w}=\{\theta+S\tau_{S,\beta}:\beta,\theta\}$ for $Q_{0}=E_{0}(Y\mid S,W,A)$, where $\theta$ is now an unspecified function of $W,A$. For clarity, we reuse the notations $\beta$ and $\theta$, although they differ from those used in the previous section for $\tilde{\Psi}$. The context, whether referring to $\tilde{\Psi}$ or $\Psi^{\#}$, will make the distinction clear. The working model ${\cal Q}_{w}$ further implies a working model ${\cal M}_{S,w}=\{P\in{\cal M}:Q_{P}\in{\cal Q}_{w}\}\subset{\cal M}$ for the observed data distribution $P_{0}$. Let $\tau_{S,\beta_{P}}\in{\cal T}_{S,w}$ be a projection of $\tau_{S,P}$ onto this working model ${\cal T}_{S,w}$ with respect to a specified loss function. We again recommend the squared-error projection of $Q_{P}$ on $\mathcal{Q}_{w}$:
$Q_{w,P}=\arg\min_{Q\in{\cal Q}_{w}}P\left(Q_{P}-Q\right)^{2}.$
Similar to Lemma 3, we have that
$\tau_{S,\beta_{P}}=\arg\min_{\tau\in\mathcal{T}_{S,w}}E_{P}\Pi_{P}(1-\Pi_{P})(1\mid A,W)(\tau_{S,P}-\tau)^{2}.$
That is, the squared-error projection of the outcome regression $Q_{P}$ onto the semi-parametric regression working model $\mathcal{Q}_{w}$ corresponds to a weighted squared-error projection of the conditional average RCT-enrollment effect $\tau_{S,P}$ onto the corresponding working model $\mathcal{T}_{S,w}$. The weights $\Pi_{P}(1-\Pi_{P})(1\mid W,A)$ stabilize the efficient influence function for $\beta_{P}$, which explains why the loss function $L_{\bar{Q},\Pi}(\tau)$ for $\tau_{S,P}$ for estimating $\tau_{S,0}$ avoids any inverse weighting by the RCT-enrollment probability, allowing for more stable and robust estimation. Just like for learning the CATE, a valid and practical loss function to learn the conditional average RCT-enrollment effect would be
$L_{\bar{Q}_{P},\Pi_{P}}(\tau)\equiv\left(Y-\bar{Q}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))\tau\right)^{2},$
We could then use a relaxed-HAL paired with this loss function with nuisance estimators $\bar{Q}_{n}$ for $\bar{Q}$ and $\Pi_{n}$ for $\Pi_{0}$. We summarize the discussion above in Lemma 6.
Lemma 6.
Let ${\cal Q}_{w}=\{\theta+S\tau_{S,\beta}:\beta,\theta\}$, where $\tau_{S,\beta}(W,A)=\sum_{j}\beta(j)\phi_{j}(W,A)$. Let $Q_{w,P}=\arg\min_{Q\in{\cal Q}_{w}}P(Q_{P}-Q)^{2}$, where $Q_{P}=E_{P}(Y\mid S,W,A)$. Let $\tau_{S,P}(W,A)=E_{P}(Y\mid S=1,W,A)-E_{P}(Y\mid S=0,W,A)$. We have that $Q_{w,P}=E_{P}(Y\mid W,A)+(S-\Pi_{P}(1\mid W,A))\sum_{j}\beta_{P}(j)\phi_{j}(W,A)$, where
$\beta_{P}=\arg\min_{\beta}E_{P}\Pi_{P}(1-\Pi_{P})(1\mid W,A)\left(\tau_{S,P}(W,A)-\sum_{j}\beta(j)\phi_{j}(W,A)\right)^{2}.$
Thus, $\sum_{j}\beta_{P}(j)\phi_{j}$ represents the projection of $\tau_{S,P}(W,A)$ onto the working model $\{\sum_{j}\beta(j)\phi_{j}:\beta\}$ with respect to a weighted $L^{2}$-norm with weights $\Pi_{P}(1-\Pi_{P})(1\mid W,A)$. We also have
$\beta_{P}=\arg\min_{\beta}E_{P}\left(Q_{P}-(S-\Pi_{P}(1\mid W,A))\sum_{j}\beta(j)\phi_{j}(W,A)\right)^{2},$
and thereby
$\beta_{P}=\arg\min_{\beta}E_{P}\left(Y-(S-\Pi_{P}(1\mid W,A))\sum_{j}\beta(j)\phi_{j}(W,A)\right)^{2}.$
Here, one could replace $Y$ by $Y-E_{P}(Y\mid W,A)$ as well. This also shows that
$\tau_{S,\beta_{P}}=\arg\min_{\tau\in\mathcal{T}_{S,w}}E_{P}(Y-E_{P}(Y\mid W,A)-(S-\Pi_{P}(1\mid W,A))\tau)^{2}.$
Therefore, we can use the following loss function for learning $\tau_{S,P}$:
$L_{\bar{Q}_{P},\Pi_{P}}(\tau)=(Y-\bar{Q}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))\tau)^{2},$
where the loss function is indexed by two nuisance parameters $\bar{Q}_{P}(W,A)=E_{P}(Y\mid W,A)$ and $\Pi_{P}(1\mid W,A)=P(S=1\mid W,A)$. The loss function is double robust with respect to misspecification of $\bar{Q}_{P}$ and $\Pi_{P}$ in the sense that $\tau_{S,\beta_{P}}=\arg\min_{\tau\in\mathcal{T}_{S,w}}E_{P}L_{\bar{Q},\Pi}(\tau)$ if either $\Pi=\Pi_{P}$ or $\bar{Q}=\bar{Q}_{P}$.
Given the definition of the semiparametric regression working model for the conditional average RCT-enrollment effect in Lemma 6, which implies a semiparametric regression working model $\mathcal{M}_{S,w}$ for $P_{0}$, we can now define a corresponding projection parameter for $\Psi^{\#}(P)$. Specifically, let $\Psi^{\#}_{{\cal M}_{S,w}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$ be defined as
$\Psi^{\#}_{{\cal M}_{S,w}}(P)\equiv E_{P}[\Pi_{P}(0\mid W,0)\tau_{S,\beta_{P}}(W,0)]-E_{P}[\Pi_{P}(0\mid W,1)\tau_{S,\beta_{P}}(W,1)].$
5.2 Canonical gradient of $\Psi^{\#}_{{\cal M}_{S,w}}$ at $P$
Let’s first find the canonical gradient of the $\beta_{P}$-component.
Lemma 7.
Let $I_{P}=E_{P}\Pi(1-\Pi)(1\mid W,A){\bf\phi}{\bf\phi}^{\top}(W,A)$. Let $\beta_{P}$ be defined as in Lemma 6 on a nonparametric model. The canonical gradient of $\beta$ at $P$ is given by:
$D_{\beta,P}=I_{P}^{-1}(S-\Pi_{P}(1\mid W,A)){\bf\phi}(W,A)\left(Y-\bar{Q}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))\sum_{j}\beta_{P}(j)\phi_{j}(W,A)\right),$
where $\bar{Q}_{P}(W,A)=E_{P}(Y\mid W,A)$, $\Pi_{P}(1\mid W,A)=P(S=1\mid W,A)$. This can also be written as
$D_{\beta,P}=I_{P}^{-1}(S-\Pi_{P}(1\mid W,A)){\bf\phi}(W,A)(Y-Q_{w,P}(S,W,A)).$
The proof can be found in Appendix A. Note that the weights $\Pi_{P}(1-\Pi_{P})$ help stabilize the variance. We then need to find the canonical gradient of the projection parameter $\Psi^{\#}_{{\cal M}_{S,w}}$ at $P$.
Lemma 8.
The canonical gradient of $\Psi^{\#}_{{\cal M}_{S,w}}$ at $P$ is given by
$D^{\#}_{{\cal M}_{S,w},P}=D^{\#}_{{\cal M}_{S,w},P_{W},P}+D^{\#}_{{\cal M}_{S,w},\Pi,P}+D^{\#}_{{\cal M}_{S,w},\beta,P},$
@@ 5.2 Canonical gradient of $\Psi^{\#}_{{\cal M}_{S,w}}$ at $P$
$D^{\#}_{{\cal M}_{S,w},P_{W},P}$ $=\Pi_{P}(0\mid W,0)\sum_{j}\beta_{P}(j)\phi_{j}(W,0)-\Pi_{P}(0\mid W,1)\sum_{j}\beta_{P}(j)\phi_{j}(W,1)-\Psi^{\#}_{{\cal M}_{S,w}}(P)$
$D^{\#}_{{\cal M}_{S,w},\Pi,P}$ =$=\left\{\frac{A}{g_{P}(1\mid W)}\tau_{S,\beta_{P}}(W,1)-\frac{1-A}{g_{P}(0\mid W)}\tau_{S,\beta_{P}}(W,0)\right\}(S-\Pi_{P}(1\mid W,A))$
$D^{\#}_{{\cal M}_{S,w},\beta,P}$ $=\sum_{j}D_{\beta,P,j}E_{P}\Pi_{P}(0\mid W,0)\phi_{j}(W,0)-\sum_{j}D_{\beta,P,j}E_{P}\Pi_{P}(0\mid W,1)\phi_{j}(W,1),$
@@
$D_{\beta,P}=I_{P}^{-1}(S-\Pi_{P}(1\mid W,A)){\bf\phi}(W,A)(Y-Q_{w,P}(S,W,A)).$
@@
5.3 TMLE of the projection parameter $\Psi^{\#}_{{\cal M}_{S,w}}$
@@
A plug-in estimator of the projection parameter $\Psi^{\#}_{{\cal M}_{S,w}}(P_{0})$ requires an estimator of $\Pi_{0}$, $\beta_{0}$, $P_{W,0}$, and a model selection method, as discussed earlier, so that the semiparametric regression working model ${\cal M}_{S,w}$ is known, including the working model $\{\tau_{S,\beta}:\beta\}$ for $\tau_{S,0}$. We then need to solve the score equation for each component of the canonical gradient of $\Psi^{\#}_{{\cal M}_{S,w}}$. First, for the $P_{W}$-component, we use the empirical distribution $P_{W,n}$ of $W$ as the estimator of $P_{W,0}$. Therefore, $P_{n}D^{\#}_{\mathcal{M}_{S,w},W,\beta,\Pi,P_{W,n}}=0$ for any $\beta,\Pi$. Then, for the $\beta$-component, we need to construct nuisance estimators $\bar{Q}_{n}$ of $\bar{Q}_{0}$ and $\Pi_{n}$ of $\Pi_{0}$. Then, with these two nuisance estimates, we can compute the MLE over the working model for $\tau_{S,0}$ with respect to our loss function $L_{\bar{Q}_{n},\Pi_{n}}(\tau)$:
$\beta_{n}^{\star}=\arg\min_{\beta}P_{n}L_{\bar{Q}_{n},\Pi_{n}}(\tau_{S,\beta}).$
@@
We have then solved $P_{n}D^{\#}_{{\cal M}_{S,w},\beta,\beta_{n}^{\star},\Pi_{n}}(O)=0$. Finally, for the $\Pi$-component, note that we have the representation: $D^{\#}_{{\cal M}_{S,w},\Pi,P}(O)=C(g,\beta)(S-\Pi_{P}(1\mid W,A))$, where we refer to $C(g,\beta)$ as the clever covariate. With an estimator $g_{n}$ of $g_{0}$, we can compute a targeted estimator $\Pi_{n}^{\star}$ solving $P_{n}C(g_{n},\beta_{n}^{\star})(S-\Pi_{n}^{\star}(1\mid W,A))=0$. The desired TMLE is then given by $\Psi^{\#}_{\mathcal{M}_{S,w}}(P_{n}^{\star})\equiv\Psi^{\#}_{{\cal M}_{S,w}}(P_{W,n},\Pi_{n}^{\star},\beta_{n}^{\star})$. For inference with respect to the projection parameter $\Psi^{\#}_{{\cal M}_{S,w}}(P_{0})$, we use that
$\Psi^{\#}_{{\cal M}_{S,w}}(P_{n}^{\star})-\Psi^{\#}_{{\cal M}_{S,w}}(P_{0})=P_{n}D^{\#}_{{\cal M}_{S,w},P_{0}}+o_{P}(n^{-1/2}),$